OpenStudy (anonymous):

use the distributive property to remove the parenthesis: (x - 12) 10 =

OpenStudy (shamil98):

(x-12)(10) multiply each value inside (x-12) to 10 to expand.

OpenStudy (anonymous):

still confused

OpenStudy (shamil98):

ex: (x + 2)(x + 5) = x·x + x·5 + 2·x + 2·5 for this you have ( x - 12 ) ( 10 + 0 ) = x * 10 + x * 0 - 12 * 10 - 12 * 0 = 10x - 120 make sense?

OpenStudy (anonymous):

ok so ultimately, i am just multiplying 10 x 12 to come up with my answer

OpenStudy (shamil98):

no no, you are multiplying everything inside the paranthese to 10 then adding the products up

OpenStudy (anonymous):

you make it look very easy but clearly this subject is not simple to me..:-)

OpenStudy (anonymous):

so what if it is an additon problem like: 10 (3 + x)? do i still add even if the ones within the parenthesis?

OpenStudy (kirbykirby):